1) The Graduate Record Examination (GRE) is a standardized test that students usually take before entering graduate school. According to the document Interpreting Your GRE Scores, a publication of the Educational Testing Service, the scores on the verbal portion of the GRE are approximately normally distributed with a mean of 462 points and a standard deviation of 119 points. Find the 90th percentile for this exam score.  Please show your work, and steps. If there is a bell curve write one.
Answer #1

Solution: We are given:

\boldsymbol{\mu=462, \sigma=119}

Let x be the 90th percentile for this exam score.

We have to find the z-value corresponding to area = 0.90. Using the standard normal table, we have:

\boldsymbol{z(0.90)=1.28 }

The excel function to find this value is:

\boldsymbol{=NORMSINV(0.90)}

Now using the z-score formula, we have:

z= \frac{x-\mu}{\sigma}

1.28= \frac{x-462}{119}

1.28 \times 119= x-462

152.32= x-462

x=462+152.32

\boldsymbol{x=614.32 }

Therefore, the 90th percentile for this exam score is 614.32

php1wQeC6.png